To investigate the effectiveness of a diet, a random sample of 16 female patients is drawn from a population of adult females using the diet. The weight of each individual in the sample is taken at the start and at the end of the diet. Assume that the population of differences in weight before and after the diet follows a normal distribution. Suppose the mean decrease in weights over all 16 subjects in the study is 4.0 pounds with the standard deviation of differences computed as 6.4 pounds. Let µx - µy = mean weight before the diet - mean weight after the diet.

In order to test if the diet is effective, what is the value of the test statistic?

Group of answer choices

2.5

2.0

1.8

1.3
Answer #1

Answer

Given that

dbar = 4 (mean difference)

sd (standard deviation for difference) = 6.4

sample size n = 16

test statistic t

t = d(savn) = 4/(6.4/V16) = 4/1.6 = 2.5

option A is correct